# ObjectFinder: Open-Vocabulary Assistive System for Interactive Object Search by Blind People

### v1

@@ Abstract.
Assistive technology can be leveraged by blind people when searching for objects in their daily lives. We created ObjectFinder, an open-vocabulary interactive object-search prototype, which combines object detection with scene description and navigation. It enables blind persons to detect and navigate to objects of their choice. Our approach used co-design for the development of the prototype. We further conducted need-finding interviews to better understand challenges in object search, followed by a study with the ObjectFinder prototype in a laboratory setting simulating a living room and an office, with eight blind users. Additionally, we compared the prototype with BeMyEyes and Lookout for object search. We found that most participants felt more independent with ObjectFinder and preferred it over the baselines when deployed on more efficient hardware, as it enhances mental mapping and allows for active target definition. Moreover, we identified factors for future directions for the development of object-search systems.
People with visual impairments, Assistive technology,
copyright: acmcopyrightjournalyear: 2025doi: XXXXXXX.XXXXXXXprice: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Human-centered computing Accessibility systems and toolsccs: Computing methodologies Computer vision tasks
@@ 1. Introduction
We all regularly visit unfamiliar buildings and shared spaces, and need to deal with uncertainty about our environment. This also includes blind people  (Müller et al., 2022; Zhang et al., 2021), for whom rapidly changing environments can be challenging to navigate, e.g., finding the way around a room, detecting unexpected obstacles (Alkhanifer and Ludi, 2015; Jeamwatthanachai et al., 2019) and searching for objects without that may have been moved by others. When first entering a new or changed environment, it is important to get an overview of the arrangement of furniture, sanitary facilities, and room installations in order to build up a mental map of a room (Façanha et al., 2020). Here, commonly leveraged haptic exploration of the environment can have safety implications due to the need for physical contact with potentially hazardous objects or surfaces (Zheng et al., 2024; Gonzalez Penuela et al., 2024), such as knifes and hot drinks. Assistive technology has attempted to fill this gap, with some applications (Eyes, 2024; Aira Company, 2024) connecting blind individuals with remote volunteers, enforcing reliance on non-disabled persons for assistance. Other approaches leverage AI  (Eyes, 2024; Achiam et al., 2023; Corporation, 2024; Lee et al., 2022a; Liu et al., 2023) to facilitate exploration of unfamiliar environments.
@@
Such technologies can be divided into two categories: description-based and detection-based systems. Description-based systems provide either a vivid and detailed descriptions of a photo (Eyes, 2024) or brief captions (Corporation, 2024; Lee et al., 2022a; Liu et al., 2023). However, these systems are not able to localize a specific object in an unfamiliar environment while ensuring that the target object is in the image area, as illustrated on the left side of Figure 1 (Challenge 1, C1). As far as detection-based systems are concerned (Google, 2024; V7, 2024; Zheng et al., 2024), current technological approaches only allow the search for specific objects (Constantinescu et al., 2020a; Yi et al., 2013; Schauerte et al., 2012; Ahmetovic et al., 2020; Kacorri et al., 2017; Google, 2024; V7, 2024). Therefore, when using current detection-based systems in an unfamiliar environment, blind users may not know what there is in a room and miss unexpected items that could be of interest, such as a plate of cookies in an unknown office, as shown on the right side of Figure 1 (Challenge 2, C2).
To address these challenges, we designed ObjectFinder, an interactive prototype for open-vocabulary object search by combining both description-based and detection-based methods. The integration of Open-Vocabulary Object Detectors (OVOD) (Cheng et al., 2024; Li et al., 2024) and Multimodal Large Language Models (MLLM) (Achiam et al., 2023; Liu et al., 2024) in ObjectFinder makes it possible to develop an interactive process ranging from object detection to description: input any target to get detailed information about the surrounding region of interest. This, in turn, identifies potentially interesting and unexpected targets, which can then be searched for in subsequent loops.
Through our work, we seek to answer the following two research questions:
  • RQ1: How can an open-vocabulary prototype integrate detection-based and description-based technologies to support object search for blind people, and what challenges are associated with its design and development?
  • RQ2: What are users’ experience and preferences regarding the open-vocabulary interactive object search system, specifically comparing the description-based and the detection-based, with the combined (ObjectFinder) systems?
We structured our work into three phases. In the first phase, we co-designed with a blind person throughout the conception and development of our system and study design. Thus, we first identified a use case that includes the following functions: target object detection (F1), localization (F2), route planning (F3), scene description (F4), and open questions (F5). Since existing description-based and detection-based systems are not tailored for object-search tasks, we developed a competitive baseline, ObjectFinder_Base, with limited detectable objects. Then, we conducted four half-hour user-in-the-loop sessions with our blind co-designer, inviting him to explore features in different scenarios (a living room and an office), leading to the refinement of ObjectFinder. In the second phase, we conducted a semi-structured interview with eight blind participants to understand their daily targets of object search with their currently used methods, and challenges they encounter in unfamiliar or changing environments. In the third phase, we conducted an exploratory evaluation of ObjectFinder inviting the same eight participants to engage with the system, and take part in a semi-structured interview afterwards. Through thematic analysis (Fereday and Muir-Cochrane, 2006), we show that the use of ObjectFinder facilitated interactive object search because it combines crucial information about localization (distance and direction) through the detection part with relationships among objects through the description part, as well with the help in navigating to the searched objects. However, we have also found that the current system is not efficient enough, is perceived too cumbersome and is not available on the cell phone, and thus cannot be integrated into daily live.
Our work makes the following key contributions: (1) We make a system contribution and introduce an open-vocabulary interactive object-search prototype that supports location of objects and provision of relevant information. (2) We provide an overview of target objects and search challenges in the daily lives of participants under consideration of the potential of assistive technology. (3) We critically reflect on the experience and effectiveness of our prototype, and outline opportunities and tensions to be considered in the future development of open-vocabulary interactive object-search systems for real-world application.
@@ 2. Related Work
In this section, we provide an overview of existing description-based and detection-based systems for scene understanding and exploration. We also discuss the integration of multimodal large language models and interaction features in assistive technologies, which form the background of our study.
2.1. Description-based and Detection-based Systems for Blind People
@@
With the advancement of Artificial Intelligence (AI), researchers are developing supplementary AI-based systems to assist blind people in exploring scenes. These systems can be classified into description-based and detection-based. The description-based systems capture the scene and describe it only once. Seeing AI (Corporation, 2024), ImageExplorer (Lee et al., 2022a) and OpenSU (Liu et al., 2023) generate brief image captions for the scenes captured by a mobile phone and enable tactile exploration of the salient objects on the touchscreen. TapTapSee (TapTapSee, Inc., nd) is an application that generates a concise phrase about the salient object in almost real-time. Be My AI, a feature of Be My Eyes (Eyes, 2024) empowered by GPT-4, delivers vivid descriptions of the scenario and allows users to ask questions. Some other works focus on the specific features of the salient object, e.g., material (Zheng et al., 2024), transparency (Zhang et al., 2021), and various hazards (Yang et al., 2018a, b). However, people with blindness should make the object within the region of interest captured by mobile devices while using description-based applications (Yang et al., 2024; Gurari et al., 2018). In addition, we found that people with blindness can barely align the photos they capture with real spatial dimensions using mobile phones during our user study which is in line with the findings of (Kacorri et al., 2017). Therefore, we implemented an object detector to identify the region of interest for the description module, ensuring a precise understanding of the area where the target is located.
@@
Detection-based systems are designed to provide real-time outputs of identified objects or features of interest. Lookout (Google, 2024) and AIPoly (V7, 2024) exemplify this capability by identifying the nearest object within the phone’s field of view. Various studies have developed wearable systems (Liu, 2023; Islam et al., 2023; Sugashini and Balakrishnan, 2024) with similar functionalities, offering real-time object information through multiple interaction modes. Research has explored the detection of personal objects using methods such as SIFT (Constantinescu et al., 2020a; Yi et al., 2013; Schauerte et al., 2012) and advanced deep learning networks (Ahmetovic et al., 2020; Kacorri et al., 2017). Navigation assistive systems utilize detection-based methods for obstacle avoidance (Bala et al., 2023; Liu et al., 2021a; Ou et al., 2022), risk assessment (Wang et al., 2024), object finding (Li et al., 2023; Duh et al., 2020; Hu et al., 2022) and passable path planning (Zou et al., 2023; Surougi and McCann, 2023; Hong et al., 2024; Jain et al., 2023). These systems autonomously select information, which may limit user agency in actively defining targets. Constantinescu et al. (Constantinescu et al., 2020b) propose a system that allows users to choose from a limited set of objects to receive audible feedback in their vicinity. In detection-based systems, the information available to users is constrained by system limitations, preventing awareness of all existing objects in the environment. This study integrates the strengths of description-based and detection-based systems to identify regions of interest and inspire users to unexpected targets.
Figure 2. Use case diagram : A blind person uses a system to find a socket in an unfamiliar office. The step with dashes is not preferred.
@@
There are eight steps in the diagram: 1) The user tells the system what he/she is looking for (e.g., a socket). 2)The system gives feedback that it understands what the user wants. 3)The system indicates it is ready for searching (through a sound). 4)The user looks around using the system. (Function 1) 5a) The system only announces sockets. 5b) The system continuously informs the user of all objects in the frame (e.g., chairs, table). 6) Socket detected. The system provides feedback with the distance, direction, and relevant points of interest (POls) in the surroundings. (Function 2, Function 4). 7a) The user wants to continue searching. The system keeps searching without intervention. 7b) The user is happy with the found object. He/she asks for further details about it, such as directions on how to reach the object and the height of the socket on the wall. (Function 5). 8) For large rooms, the system should navigate the user (e.g., ”Walk around the table on the left side”). (Function 3)
Figure 3. Hardware design and components of the wearable system ObjectFinder. It incorporates a stereo camera to capture visual information about the user’s surroundings, a pair of buttons and a microphone to collect the user’s commands. Simultaneously, it executes algorithms through a lightweight processor. To provide a comprehensive and immersive experience, the system delivers spatial-aware informational feedback directly to the user via bone-conduction headphones.

### v2

# ObjectFinder: An Open-Vocabulary Assistive System for Interactive Object Search by Blind People

@@ Abstract.
Searching for objects in unfamiliar scenarios is a challenging task for blind people. It involves specifying the target object, detecting it, and then gathering detailed information according to the user’s intent. However, existing description- and detection-based assistive technologies do not sufficiently support the multifaceted nature of interactive object search tasks. We present ObjectFinder, an open-vocabulary wearable assistive system for interactive object search by blind people. ObjectFinder allows users to query target objects using flexible wording. Once the target object is detected, it provides egocentric localization information in real-time, including distance and direction. Users can then initiate different branches to gather detailed information based on their intent towards the target object, such as navigating to it or perceiving its surroundings. ObjectFinder is powered by a seamless combination of open-vocabulary models, namely an open-vocabulary object detector and a multimodal large language model. The ObjectFinder design concept and its development were carried out in collaboration with a blind co-designer. To evaluate ObjectFinder, we conducted an exploratory user study with eight blind participants. We compared ObjectFinder to BeMyAI and Google Lookout, popular description- and detection-based assistive applications. Our findings indicate that most participants felt more independent with ObjectFinder and preferred it for object search, as it enhanced scene context gathering and navigation, and allowed for active target identification. Finally, we discuss the implications for future assistive systems to support interactive object search.
copyright: acmcopyrightjournalyear: 2025doi: XXXXXXX.XXXXXXXprice: 15.00isbn: 978-1-4503-XXXX-X/18/06
@@ 1. Introduction
Table 1. Comparison of different systems that can be used for object search.
System Purpose Enabling Source Interaction Device
RSA (Eyes, 2024; TapTapSee, Inc., nd; Aira Company, 2024) Multi-Purpose Human Dialogue Smartphone
ProgramAlly (Herskovitz et al., 2024) Object Search AI Filter Customization Smartphone
WorldScribe (Chang et al., 2024) Exploration AI Intent Customization Smartphone
WanderGuide (Kuribayashi et al., 2025) Exploration AI Dialogue, Button-Driven Option Selection Robot
BeMyAI (Eyes, 2023) Description AI Dialogue Smartphone
Lookout (Google, 2024) Exploration AI - Smartphone
Find My Things (Wen et al., 2024) Object Search AI Teachable Object Recognition Smartphone
LifeInsight (Mathis and Schöning, 2025) Question and Answering AI Dialogue Wearable Device
ObjectFinder Object Search & Exploration AI Dialogue, Button-Driven Option Selection Wearable Device
Blind people often face challenges when searching for objects in unfamiliar environments (Müller et al., 2022; Zhang et al., 2021). Independently searching for a specific object, such as locating the nearest available chair in a spacious lobby through haptic exploration, can be particularly difficult. To search for an object, users would first query the target object, and detect a candidate. In order to determine if it is the desired one, both egocentric (e.g. 11 o’clock, 2.4 meters away”) and allocentric (e.g.next to the desk”) information are necessary for a blind user to perceive the object in their environment (Golledge, 1999; Martolini et al., 2021). Upon locating the target object, the user’s intent may vary. As shown in examples of Figure 1, once the coffee table is detected, the user might prefer an audio description of items on the table over immediate physical interaction. Conversely, if the target is a fan, the user might wish to navigate towards it to turn it on. Throughout the search, there may also be a desire to explore the surroundings for better navigation and potentially discover other targets for subsequent searches (Jain et al., 2023b).
@@
Such an interactive object search task is multifaced, however, no current AI-powered assistive technology can yet handle all the associated subtasks. We categorize the existing assistive technologies for blind people into description-based and detection-based systems. Description-based systems provide vivid and detailed descriptions of a photo (Eyes, 2023) or brief captions (Corporation, 2024; Lee et al., 2022a; Liu et al., 2023). However, these systems are unable to localize a specific object in an unfamiliar environment while ensuring that the target object is in the frame (Xie et al., 2024) (Challenge 1, C1). Detection-based systems (Google, 2024; V7, 2024; Zheng et al., 2024), on the other hand, either allow only the search for a limited number of pre-defined objects (Constantinescu et al., 2020a; Yi et al., 2013; Schauerte et al., 2012; Ahmetovic et al., 2020; Kacorri et al., 2017; Google, 2024; V7, 2024; Wen et al., 2024) or provide filtered information (Herskovitz et al., 2024; Chang et al., 2024), limiting understanding of unfamiliar scenes. Therefore, when using current detection-based systems in an unfamiliar environment, blind users may not know what is in a room comprehensively and miss items that could be of interest (Challenge 2, C2). Apart from that, both detection-based and description-based systems fail to provide egocentric information (distance and direction) or support question and answering directed towards the target.
The challenges also exist in remote sighted assistance (RSA). The procedure by which the remote agents (Aira Company, 2024; Eyes, 2024; TapTapSee, Inc., nd) help to identify objects and describe surroundings involves capturing images from the video feed and zooming in to obtain the necessary visual information (Lee et al., 2020). In this context, it is time-consuming for remote agents to adjust the video frame, and they find it challenging to continuously orient the users (Xie et al., 2022a). Moreover, recognizing landmarks presents significant difficulties for the agents (Lee et al., 2022b; Xie et al., 2022b; Kamikubo et al., 2020). Thus in this work, we aim to address the following question:
How to integrate the advantages of description- and detection-based assistive systems to support interactive object search by blind people?
To this end, we designed ObjectFinder, which seamlessly combines open-vocabulary models, an Open-Vocabulary Object Detector (OVOD)(Cheng et al., 2024) and a Multimodal Large Language Model (MLLM)(Achiam et al., 2023), to facilitate an interactive process that ranges from object detection to description for object search. Users can input any target via voice commands for object detection, then scan the scene. Once a candidate is detected, the system will notify the user to stand still to orient to the target and output real-time egocentric information (distance and direction). Following this, users can acquire comprehensive information about their surroundings, tailored to their intent, based on the keyframe captured at the time of detection. This process facilitates the identification of potentially interesting and unexpected targets, which can then be explored in further detail during subsequent iterations.
Based on prior works in object search tasks, we formulated design considerations for an interactive system that enables flexible querying, supports various search subtasks, and adapts the system feedback to user intent. We co-designed with a blind person throughout the conception and development of ObjectFinder, across two months and four iterations. We deconstruct the complex object search task into the following functions: target object detection, localization, route planning, scene description, and open questions. The pipeline for integrating all functions and interaction features was refined based on the iterative feedback from the blind co-designer.
To evaluate the effectiveness and efficiency of ObjectFinder in facilitating object search, we conducted an exploratory evaluation with eight blind participants. They engaged with the system and participated in a semi-structured interview afterwards. BeMyAI (Eyes, 2023) and Google Lookout (Google, 2024), popular commercial description- and detection-based systems, were used as baseline comparisons. Through thematic analysis (Fereday and Muir-Cochrane, 2006), we demonstrate that the use of ObjectFinder enhanced interactive object search. It integrated crucial information about egocentric localization (distance and direction) from the detection component, and allocentric relationships among objects from the description component. Additionally, route planning was a valuable feature of ObjectFinder for searching objects. Although ObjectFinder provides feedback based on users’ intents, individual variations in procedures and information preferences, also influenced by a scenario’s scope and familiarity, underscore the need for customization and personalization, as discussed later. ObjectFinder represents a significant advancement in bridging the technological gap in object search, particularly in unfamiliar contexts. Taking advantage of both description- and detection-based systems, its technical approach has the potential to find systems broadly to enhance the independence of blind individuals in their daily lives.
@@ 2. Related Work
In this section, we introduce the task of object search for blind individuals and provide an overview of existing description-based and detection-based systems designed specifically for them, which can partially address the task. Since no assistive system currently exists for object search in unfamiliar scenarios, we refer to procedures from embodied AI for object search, which typically mimic human behavior. This forms the background knowledge for our study.
2.1. Object Search in Unfamiliar Scenarios
Object search is a multifaceted task that involves object detection, exploration, navigation, and more. In addition to small items that blind individuals frequently search for in their daily lives such as smartphone, keys and wallet (netz-barrierefrei.de, nd), they often search for large, salient objects as landmarks to improve orientation in unfamiliar environments (Project, nd; Yang et al., 2010). When searching for objects in unfamiliar environments, blind people typically seek an initial overview of the space, followed by specific details as required (Chang et al., 2024; Shneiderman, 1996). If the target object has been found, blind people may have various intentions regarding it. For example, they might navigate to the object to interact with it (Herskovitz et al., 2023) (e.g., find a free chair and sit on it), identify a specific object (Brady et al., 2013; Hong and Kacorri, 2024) (e.g., check whether a bottle is shampoo), or perceive the surroundings of the object (e.g., the tabletop (Herskovitz et al., 2023)), which may be too far or inconvenient to touch (Gonzalez Penuela et al., 2024). A participant in (Herskovitz et al., 2023) defined the use case of locating an empty chair in the classroom and imagined how the object search system should work: he preferred to scan the environment with smart glasses rather than waving his phone in the crowd, then the system find an empty chair and give directions on how to walk to the chair.
Some technologies have been proposed to partially address the challenges of object search (Table 1). Vizwiz-LocateIt (Bigham et al., 2010) lets users photograph target objects, ask questions to a remote worker on Mechanical Turk, and navigate via sonification. Tools such as AIRA (Aira Company, 2024), Vizwiz (Gurari et al., 2018), and BeMyEyes (Eyes, 2024) utilize crowdsourcing to connect blind people with sighted people for real-time remote sighted assistance including object search. However, asking the blind people to move their phone to adjust the video frame is time-consuming (Lee et al., 2020). WanderGuide (Kuribayashi et al., 2025) has subfunctions for object search implemented on a suitcase, but is designed primarily for exploration without specific consideration of the object search procedure. Bhanuka et al. (Gamage et al., 2023) suggest that the conversational interface on wearable devices is suitable for the complex task of providing environmental information. We categorize existing AI-based assistive technology related to object search into description-based and detection-based systems.
2.2. Description-based Systems for Blind People
@@
The description-based systems capture the scene and describe it only once. Seeing AI (Corporation, 2024), ImageExplorer (Lee et al., 2022a), and OpenSU (Liu et al., 2023) generate brief image captions for the scenes captured by a mobile phone and enable tactile exploration of the salient objects on the touchscreen. TapTapSee (TapTapSee, Inc., nd) is an application that generates a concise phrase about the salient object in almost real-time. BeMyAI (Eyes, 2023), a feature of BeMyEyes empowered by GPT-4, delivers vivid descriptions of the scenario and allows users to ask questions. Research on BeMyAI (Xie et al., 2024, 2025) indicates that while it serves as a form of distributed cognition, it faces challenges in intent recognition and frequently necessitates the use of multiple images to accurately convey information. NaviGPT (Zhang et al., 2025) is a mobile navigation system that provides a brief description of the road ahead. LifeInsight (Mathis and Schöning, 2025) is a wearable system embedded with GPT-4 for question answering. Some other works focus on the specific features of the salient object, e.g., material (Zheng et al., 2024), transparency (Zhang et al., 2021), and various hazards (Yang et al., 2018a, b). However, people with blindness should make the object within the region of interest captured by mobile devices while using description-based applications (Yang et al., 2024; Gurari et al., 2018). In addition, we found that people with blindness can barely align the photos they capture with real spatial dimensions using mobile phones during our user study, which is in line with the findings of (Kacorri et al., 2017). However, Gonzales et al. (Gonzalez Penuela et al., 2024) determined that the primary goal for users of AI-powered scene description applications is to identify specific objects. Therefore, we implemented an object detector to identify the region of interest for the description module, ensuring a precise understanding of the area where the target is located.
2.3. Detection-based Systems for Blind People
@@
Detection-based systems are designed to provide real-time outputs of identified objects or features of interest. Lookout (Google, 2024) and AIPoly (V7, 2024) exemplify this capability by identifying the nearest object within the phone’s field of view. Various studies have developed wearable systems (Liu, 2023; Islam et al., 2023; Sugashini and Balakrishnan, 2024) with similar functionalities, offering real-time object information through multiple interaction modes. WorldScribe (Chang et al., 2024) delivers real-time descriptions of the current view, tailoring the information based on distance and user intent. Research has explored the detection of personal objects using methods such as SIFT (Constantinescu et al., 2020a; Yi et al., 2013; Schauerte et al., 2012) and advanced deep learning networks (Ahmetovic et al., 2020; Kacorri et al., 2017; Wen et al., 2024). ProgramAlly (Herskovitz et al., 2024) allows users to customize the information filter and efficiently detect specific features of an object. This suggests that a predefined list is not preferred for exploration. Navigation assistive systems utilize detection-based methods for obstacle avoidance (Bala et al., 2023; Liu et al., 2021; Ou et al., 2022), risk assessment (Wang et al., 2024), object finding (Li et al., 2023; Duh et al., 2020; Hu et al., 2022), shopping (Boldu et al., 2020) and passable path planning (Zou et al., 2023; Surougi and McCann, 2023; Hong et al., 2024; Jain et al., 2023a). These systems autonomously select information, which may limit user agency in actively specifying targets. Constantinescu et al. (Constantinescu et al., 2020b) propose a system that allows users to choose from a limited set of objects to receive audible feedback in their vicinity. In detection-based systems, the information available to users is limited by system constraints, which prevents them from gaining a comprehensive understanding of essential scene context for object search.
2.4. Reference Procedure for Object Search
Since no existing AI system fully addresses the challenges of object search for blind users, we examine the object-search procedures of embodied AI to guide the design of an assistive object search system. Object search is widely recognized as a challenging task that integrates both perceptual and cognitive processes (Sun et al., 2025). Typically, embodied agents (Aydemir et al., 2013; Chaplot et al., 2020; Yokoyama et al., 2024) first receive an object query from the user, analyze their surroundings, hypothesize the potential location of the target object, and then plan a navigation path accordingly. Recent workflows leveraging LLMs, such as UniGoal (Yin et al., 2025) and SG-Nav (Yin et al., 2024), allow robots to continuously explore their environment and match discovered objects with the intended target. CogNav (Cao et al., 2024) investigates the modeling of cognitive process of object search, which involves a broad and contextual search back and forth to build a cognitive map. Upon observing the target, it verifies the candidate according to the vicinity, then confirms the candidate. Taioli et al. (Taioli et al., 2024) equipped the object search agent with a self-questioner and an interaction trigger to produce a refined detection description that includes dialogue regarding the target object. In this study, we explore the subtasks involved in object search and integrate them into a unified pipeline specifically designed for blind users, taking advantage of both description- and detection-based systems.
Figure 2. Martin walks into an unfamiliar office and uses an object-search system to search for a socket to charge his smartphone. (a) Martin first specifies the target to the system, which then repeats it for confirmation. (b) While scanning, candidates are detected. The socket “4 meters away at his 10 o’clock next to the trash bin” is not what he wants. (c) However, another socket “3 meters away at his 2 o’clock next to the workstation” is the desired one, as he plans to study there. (d) After confirming the target, Martin may ask for more details. In large rooms, the system should navigate him to the socket.
@@ 2.4. Reference Procedure for Object Search
Martin walks into an unfamiliar office and uses an object-search system to search for a socket to charge his smartphone. (a) Martin first specifies the target to the system, which then repeats it for confirmation. (b) While scanning, candidates are detected. The socket “4 meters away at his 10 o’clock next to the trash bin” is not what he wants. (c) However, another socket “3 meters away at his 2 o’clock next to the workstation” is the desired one, as he plans to study there. (d) After confirming the target, Martin may ask for more details. In large rooms, the system should navigate him to the socket.
3. ObjectFinder
ObjectFinder is a wearable prototype designed for interactive object search. Blind users can specify their target using flexible wording. Once the target is detected, they receive real-time egocentric localization information. They can further obtain detailed feedback based on their intentions toward the targets. We co-designed ObjectFinder with a blind person P0 (see Table 2) by proposing an envisioned scenario, constructing an initial prototype based on it, and then conducting four refinement iterations over two months.
3.1. Design Goals
Overall, drawing on related works in object search, we designed ObjectFinder with three primary goals:
Providing flexibility in target queries and information retrieval. Blind people prefer to query flexible target objects and their surroundings, often discovering new items of interest. ObjectFinder should facilitate seamless conversational interactions using open-vocabulary models to bridge wording gaps.
Supporting various subtasks during search. Object search is a complex task involving several sequential subtasks: target specification, detection, and feedback generation. ObjectFinder should simplify this process by organizing these subtasks into a user-friendly pipeline with accessible interaction features.
Adapting to the user intent of the target object. Blind people exhibit varying intents for identified targets, from navigation to mere perception, and require tailored descriptions or guidance based on the familiarity of their environment. ObjectFinder should offer options that allow users to gather system feedback based on their specific intents.
Figure 3. Hardware design and components of the wearable system ObjectFinder. It incorporates a stereo camera to capture visual information about the user’s surroundings, a pair of buttons, and a microphone to collect the user’s commands. Simultaneously, it executes algorithms through a lightweight processor. To provide a comprehensive and immersive experience, the system delivers spatial-aware informational feedback directly to the user via bone-conduction headphones.